# Screen Reader Users in the Vibe Coding Era: Adaptation, Empowerment, and New Accessibility Landscape

### v1

@@ 7. Conclusion
Table A shows the think-aloud questions asked in the initial interview.
Table A. Question phrases by category for the initial interview.
@@
Background How long have you been learning programming?
(Pre-task) Can you share your work experience?
Have you used AI coding assistants before (e.g., GitHub Copilot, ChatGPT)? Which ones?
How frequently do you use them?
What kind of tasks do you typically use them for? What benefits do you think they provide?
Thinking What was your overall thought process when completing the task?
Interaction Did you find learning and mastering GitHub Copilot easy or difficult? What aspects influenced this experience?
Can you describe how easy or difficult it is to understand Copilot’s suggestions?
How do you decide which suggestions to accept or reject? How do you evaluate Copilot’s suggestions?
How much do you trust Copilot? What could Copilot do to make you trust it more?
Have you encountered situations where Copilot generated incorrect answers? How did you handle such situations?
Among Copilot’s different features (code completion, inline chat, ask, edit, agent mode), which do you prefer and why? Does it vary by situation? Any specific examples? Why didn’t you use certain modes?
Copilot’s Impact In your daily programming activities, which tasks or stages do you find most challenging? Please provide examples.
Does GitHub Copilot help alleviate the programming challenges you mentioned? Please discuss with specific examples.
Additionally, for which tasks do you think Copilot provides the most help?
Difficulties & What difficulties or challenges have you encountered while using Copilot? If any, what were they?
Expectations Would you consider using GitHub Copilot long-term?
What additional features or improvements would you like Copilot to provide to help you more?
Table B shows the think-aloud questions asked in the follow-up interview.
Table B. Question phrases by category for the follow-up interview.
@@
Frequency of Use How often have you used Copilot in the past two weeks? Why?
Modes Which modes or features did you use more frequently during these two weeks?
What specific tasks did you use them for?
What benefits or shortcomings did you notice in their usage?
Usage Do you think Copilot understands your needs well?
Experience How do you prefer to review responses now?
& Impact Do you find the process of reviewing and accepting responses smooth? Were there any obstacles or inconveniences?
After continuous use of Copilot, has your overall impression of the tool changed? What new insights or discoveries have you made?
Challenges & During the past two weeks, did you encounter any new difficulties or challenges while using Copilot?
Feedback After using Copilot for a while, what features or improvements would you most like to see?
Table C shows the questions rated on a 1-7 Likert scale in both the initial and follow-up interviews, where 1 indicates strong disagreement and 7 indicates strong agreement.
Table C. Likert scale questions’ IDs and corresponding phrases.
QID Phrase
Q1 Copilot’s code suggestions make me more efficient in completing programming tasks.
Q2 Copilot helps me improve my programming skills and inspires new ideas.
Q3 I can predict when Copilot will generate useful results.
Q4 I trust Copilot’s code suggestions and am willing to use them directly.
Q5 Copilot’s suggestions won’t lead me to develop bad coding habits, such as accepting code without fully understanding it.
Q6 I can efficiently interact with Copilot via keyboard shortcuts or other methods.
Q7 Copilot’s suggestions are easy to understand and review using a screen reader.
Q8 I feel that I am actively leading the collaboration with Copilot, instead of passively following its suggestions.
Q9 Code completion distracts me and makes me feel troubled while programming.
Q10 I prefer Copilot to provide suggestions automatically rather than manually triggering them.
@@
Finally, we presented all 16 users’ interaction timeline during their programming task in Figure B.
Figure B. All users’ action timelines during their programming task.

### v2

@@ 6. Conclusion
Table C shows the semi-structured questions asked in the initial interview.
Table D shows the semi-structured questions asked in the follow-up interview.
Table B shows the questions rated on a 1-7 Likert scale in both the initial and follow-up interviews, where 1 indicates strong disagreement and 7 indicates strong agreement.
Table B. Likert scale questions’ IDs, labels and corresponding phrases.:
QID Label Phrase
Q1 Efficiency Copilot’s code suggestions make me more efficient in completing programming tasks.
Q2 Learning Copilot helps me improve my programming skills and inspires new ideas.
Q3 Predictability I can predict when Copilot will generate useful results.
Q4 Trust I trust Copilot’s code suggestions and am willing to use them directly.
Q5 Resists Overreliance Copilot’s suggestions won’t lead me to develop bad coding habits, such as accepting code without fully understanding it.
Q6 Interaction Efficiency I can efficiently interact with Copilot via keyboard shortcuts or other methods.
Q7 Easily Reviewed Copilot’s suggestions are easy to understand and review using a screen reader.
Q8 User Control I feel that I am actively leading the collaboration with Copilot, instead of passively following its suggestions.
Q9 Automation Preference I prefer Copilot to provide suggestions automatically rather than manually triggering them.
Q10 Completion Distraction Code completion distracts me and makes me feel troubled while programming.
Table C. Question phrases by category for the initial interview.:
@@
Background (Pre-task) $\cdot$ How long have you been learning programming?
$\cdot$ Can you share your work experience?
$\cdot$ Have you used AI coding assistants before (e.g., GitHub Copilot, ChatGPT)? Which ones?
$\cdot$ How frequently do you use them?
$\cdot$ What kind of tasks do you typically use them for? What benefits do you think they provide?
Thinking $\cdot$ What was your overall thought process when completing the task?
Interaction $\cdot$ Did you find learning and mastering GitHub Copilot easy or difficult? What aspects influenced this experience?
$\cdot$ Can you describe how easy or difficult it is to understand Copilot’s suggestions?
$\cdot$ How do you decide which suggestions to accept or reject? How do you evaluate Copilot’s suggestions?
$\cdot$ How much do you trust Copilot? What could Copilot do to make you trust it more?
$\cdot$ Have you encountered situations where Copilot generated incorrect answers? How did you handle such situations?
$\cdot$ Among Copilot’s different features (code completion, inline chat, ask, edit, agent mode), which do you prefer and why? Does it vary by situation? Any specific examples? Why didn’t you use certain modes?
Copilot’s Impact $\cdot$ In your daily programming activities, which tasks or stages do you find most challenging? Please provide examples.
$\cdot$ Does GitHub Copilot help alleviate the programming challenges you mentioned? Please discuss with specific examples.
$\cdot$ Additionally, for which tasks do you think Copilot provides the most help?
Difficulties & Expectations $\cdot$ What difficulties or challenges have you encountered while using Copilot? If any, what were they?
$\cdot$ Would you consider using GitHub Copilot long-term?
$\cdot$ What additional features or improvements would you like Copilot to provide to help you more?
Table D. Question phrases by category for the follow-up interview.:
@@
Frequency of Use $\cdot$ How often have you used Copilot in the past two weeks? Why?
Modes $\cdot$ Which modes or features did you use more frequently during these two weeks?
$\cdot$ What specific tasks did you use them for?
$\cdot$ What benefits or shortcomings did you notice in their usage?
Usage Experience & Impact $\cdot$ Do you think Copilot understands your needs well?
$\cdot$ How do you prefer to review responses now?
$\cdot$ Do you find the process of reviewing and accepting responses smooth? Were there any obstacles or inconveniences?
$\cdot$ After continuous use of Copilot, has your overall impression of the tool changed? What new insights or discoveries have you made?
Challenges & Feedback $\cdot$ During the past two weeks, did you encounter any new difficulties or challenges while using Copilot?
$\cdot$ After using Copilot for a while, what features or improvements would you most like to see?
@@
Finally, we presented all 16 users’ interaction timelines during their programming task in Figure A.
Figure A. All users’ action timelines during their programming task.: